# Are Graph Neural Networks Optimal Approximation Algorithms?

### v5

@@ 3.2 Message Passing for Max-CSPs
We construct our message passing algorithm as follows. We formulate SDP 1, a general SDP for Max-CSPs that is equivalent to the UGC-optimal SDP of Raghavendra (2008). Concretely, let $\text{SDP}(\Lambda)$ be the optimal value of the SDP 1 on instance $\Lambda$. Then, the approximation ratio for the Max-k-CSP problem achieved by the SDP 1 is
$\text{Approximation Ratio}\vcentcolon=\min_{\Lambda\in\text{Max-k-CSP}}\frac{\text{OPT}(\Lambda)}{\text{SDP}(\Lambda)},$
where the minimization is taken over all instances $\Lambda$ with arity $k$. Note the approximation ratio is always smaller than one. Then, there is no polynomial time algorithm that can achieve a superior (larger) approximation ratio assuming the truth of the UGC. Next, based on SDP 1 we form a quadratically penalized Lagrangian $\mathcal{L}_{p}(\mathbf{v})=\mathcal{L}_{o}(\mathbf{v})+\rho\mathcal{L}_{c}(\mathbf{v})^{2}$. We then show that gradient descent on this Lagrangian takes the form of a message-passing algorithm that can provably converge to an approximate global optimum for the SDP. A simplified version of the algorithm is provided in algorithm 1. To emphasize the conceptual point, we abuse notation in algorithm 1 and use $\mathcal{L}_{o}(v_{i},v_{j}),\mathcal{L}_{c}(v_{i},v_{j})$ to denote the terms involving $v_{i}$ and its neighbors $v_{j}$ in the objective and the constraints of the Lagrangian respectively. See equation 63 and algorithm 2 for the detailed version of the Max-CSP message passing update. See appendix A and B for Min-Vertex-Cover and Max-3-SAT message passing updates.
Our main theorem is the following.
@@
For the formal theorem and proof see Theorem C.1.
3.3 OptGNN for Max-CSP
The message-passing form of algorithm 1 allows us to define OptGNN, a natural GNN generalization that captures this gradient iteration.
Definition (OptGNN for Max-CSP).
Let $\Lambda$ be a Max-CSP instance on a constraint graph $G_{\Lambda}$ with $N$ nodes. Let $U$ be an input matrix of dimension $r\times N$ for the $N$ nodes with embedding dimension $r$. Let $\mathcal{L}_{\rho}$ be the loss defined as in equation 63 associated with the Max-CSP instance $\Lambda$. Let $M$ be the OptGNN weights which are a set of matrices $M\vcentcolon=\{M_{1},M_{2},...,M_{T}\}\in\mathbb{R}^{r\times 2r}$. We define $\text{OptGNN}_{(M,\Lambda)}:\mathbb{R}^{r\times N}\rightarrow\mathbb{R}$ to be the function
$\text{OptGNN}_{(M,\Lambda)}(U)=$ (13)
$\mathcal{L}_{\rho}\circ\text{LAYER}_{M_{T}}\circ\dots\circ\text{LAYER}_{M_{2}}\circ\text{LAYER}_{M_{1}}(U).$
Here we define $\text{LAYER}_{M_{i}}:\mathbb{R}^{r\times N}\rightarrow\mathbb{R}^{r\times N}$ to be the function $\text{LAYER}_{M_{i}}(U)=M_{i}(\text{AGG}(U))$, where $\text{AGG}:\mathbb{R}^{r\times N}\rightarrow\mathbb{R}^{2r\times N}$ is the aggregation function:
$\text{AGG}(U)\vcentcolon=\begin{bmatrix}U\\ \nabla\mathcal{L}_{\rho}(U)\end{bmatrix}.$
To update the parameters of OptGNN, We feed the output of the final layer $\text{LAYER}_{M_{T}}$ to the Lagrangian $\mathcal{L}_{p}:\mathbb{R}^{r\times N}\rightarrow\mathbb{R}$ and backpropagate its derivatives. It is important to emphasize that the data is the instance $\Lambda$ and not the collection of vectors $U$ which can be chosen entirely at random. The form of the gradient $\nabla\mathcal{L}_{\rho}$ is a message passing algorithm over the nodes of the constraint graph $G_{\Lambda}$. Therefore, OptGNN is a message passing GNN over $G_{\Lambda}$. This point is of practical importance as it is what informs out implementation of the OptGNN architecture. To make the theoretical connection between OptGNN and SDP solving precise, we must ensure that a small amount of additive stochastic noise $\{b_{t}\}_{t\in[T]}$ is added at each layer for the sake of theoretical convergence Jin et al. (2017) as $\mathcal{L}_{\rho}$ has exponential saddle points. We then arrive at the following corollary.
Corollary 1.
Given a Max-k-CSP instance $\Lambda$, there is an OptGNN${}_{(M,\Lambda)}$ with $T=poly(\log(\delta^{-1}),\epsilon^{-1},|\mathcal{P}|q^{k})$ layers, and embeddings of dimension $r=|\mathcal{P}|q^{k}$ such that there is an instantiation of learnable parameters $M=\{M_{t}\}_{t\in[T]}$ and random perturbations $b=\{b_{t}\}_{t\in[T]}$ that outputs a set of vectors $\mathbb{v}$ satisfying the constraints of SDP 1 and approximating its optimum, $\text{SDP}(\Lambda)$, to error $\epsilon$ with probability $1-\delta$ over the randomness in $b$.
Moving on from our result on representing approximation algorithms, it is natural to ask whether OptGNN is even learnable. That is to say, does OptGNN approximate the value of $\text{SDP}(\Lambda)$ when given a polynomial amount of data? We provide a perturbation analysis to establish the polynomial sample complexity of PAC-learning OptGNN. The key idea is to bound the smoothness of the polynomial circuit AGG used in the OptGNN layer. We state the informal version below. For the formal version see Lemma E.5
Lemma 3.1 (PAC learning (informal)).
Let $\Lambda$ be a dataset of Max-k-CSP instances over an alphabet of size $[q]$. Here the dataset $\Lambda\vcentcolon=\Lambda_{1},\Lambda_{2},...,\Lambda_{\Gamma}\sim\mathcal{D}$ is drawn i.i.d from a distribution over instances $\mathcal{D}$ with no more than $|\mathcal{P}|$ predicates. Let $M$ be a set of parameters $M=\{M_{1},M_{2},...,M_{T}\}$ in a parameter space $\Theta$. Then for $T=O(|\mathcal{P}|^{4}q^{4k})$, for $\Gamma=O(\frac{1}{\epsilon^{4}}|\mathcal{P}|^{6}q^{6k}\log^{4}(\delta^{-1}))$, let $\hat{M}$ be the empirical loss minimizer for arbitrary choice of initial embeddings $U$.
$\hat{M}\vcentcolon=\operatorname*{arg\,min}_{M\in\Theta}\frac{1}{\Gamma}\sum_{i\in[\Gamma]}\textup{\text{OptGNN}}_{(M,\Lambda_{i})}(U)$
Let the empirical loss be denoted
$\text{EMP}(\Lambda)\vcentcolon=\frac{1}{\Gamma}\sum_{i\in[\Gamma]}\textup{\text{OptGNN}}_{(\hat{M},\Lambda_{i})}(U)$
Then we have that OptGNN is $(\Gamma,\epsilon,\delta)$-PAC learnable. That is to say the empirical loss is within $\epsilon$ from the optimal distributional loss with probability greater than $1-\delta$:
$\Pr\left[\left|\textup{\text{EMP}}(\Lambda)-\mathbb{E}_{\Lambda\sim\mathcal{D}}\left[\textup{\text{SDP}}(\Lambda)\right]\right|\leq\epsilon\right]\geq 1-\delta.$
We note that this style of perturbation analysis is akin to the VC theory on neural networks adapted to our unsupervised setting. As such, our analysis does not provide an explanation for why deep neural networks can learn with fewer samples than parameters which is a notorious challenge in learning theory Bartlett et al. (2021). Nevertheless, our analysis sheds light on how architectures that capture gradient iterations can successfully generalize.
Algorithm 1 Message Passing for Max-CSP (informal).
  Input: Max-CSP Instance $\Lambda$
  $\rho$: penalty for the constraint}
  $\mathcal{L}_{o}$: objective component of the penalized objective}
  $\mathcal{L}_{c}$: constraint component of the penalized objective}
  $\mathbf{v}^{0}\leftarrow\text{Uniform}(\mathcal{S}^{n-1})$Initialize vectors to uniform on the unit sphere}
  for $t=1,2,\dots,T$ do
     what
     for $v_{w}^{t}\in\mathbf{v}^{t}$message passing on the constraint graph} do
        what
$\frac{\partial\mathcal{L}(\mathbf{v}^{t})}{\partial v_{i}}$ $=\sum_{j\in N(i)}\frac{\partial\left(\mathcal{L}_{\text{o}}(v_{i},v_{j})+\rho\mathcal{L}_{\text{c}}(v_{i},v_{j})^{2}\right)}{\partial v_{i}}$
${v}_{i}^{t+1}$ $\leftarrow v_{i}^{t}-\eta\frac{\partial\mathcal{L}(\mathbf{v})}{\partial v_{i}}$
        if $\|v_{i}^{t+1}-v_{i}^{t}\|\leq\psi$ then
           $v_{i}^{t+1}\leftarrow v_{i}^{t+1}+\mathcal{N}(0,\sigma I)$
        end if
     end for
  end for
  return $\mathbf{v}^{t}$
  Output: vector solution for SDP 1
@@ 3.3 OptGNN for Max-CSP
Obtaining neural certificates. The goal of OptGNN is to find high-quality solutions to CO problems by capturing powerful convex relaxations. To underscore this point, we construct dual certificates of optimality from the embeddings of OptGNN. The certificates provide a bound on the optimal solution of the SDP and can be computed directly from the embeddings produced by OptGNN. The key idea is to estimate the dual variables of SDP 1 from the output embeddings/SDP solution $\mathbf{v}$. Since we use a quadratic penalty for constraints, the natural dual estimate is one step of the augmented method of Lagrange multipliers on the SDP solution which can be obtained straightforwardly from the overparameterized primal problem variables $\mathbf{v}$. Of course, this estimate need not be dual feasible. To handle this problem we analytically bound the error in satisfying the KKT conditions of SDP 1. See C.2 for derivations and extended discussion, and 4.2 for an experimental demonstration.
3.4 OptGNN in practice
Figure 1 summarizes the OptGNN pipeline for solving CO problems. To run the forward pass, the first step is message passing. Given a graph $G$ with $N$ nodes, from some initial features as input, OptGNN computes the node embeddings $\mathbf{v}\in\mathbb{R}^{N\times r}$ using message passing per equation 13. Please refer to the pseudocode for Max-Cut in 3 and for general SDPs in 4.
After message passing, we compute $\mathcal{L}(\mathbf{v};G)$. For Max-Cut, this is $-\sum\nolimits_{(i,j)\in E}\tfrac{1}{2}(1-\langle v_{i},v_{j}\rangle)$. In training, we treat this as the loss and call the backward pass to calculate gradients. We use Adam (Kingma & Ba, 2014) to update the parameter matrices $M$. The same approach is used for other problems; see Appendix A for examples of other loss functions. Given a training distribution $\mathcal{D}$, the network is trained in a completely unsupervised fashion by minimizing $\mathbb{E}_{G\sim\mathcal{D}}[\mathcal{L}(\mathbf{v};G)]$ with a standard automatic differentiation package like PyTorch (Paszke et al., 2019).
A practical benefit of our approach is that users do not need to reimplement the network to handle each new problem. Users need only implement the appropriate Lagrangian, and our implementation uses automatic differentiation to compute the messages in the forward pass.
To obtain a discrete solution at inference time, the vectors solving the SDP must be rounded. To do this, we select a random hyperplane vector $y\in\mathbb{R}^{r}$, and for each node with embedding vector $v_{i}$, we calculate its discrete assignment $x_{i}\in\{-1,1\}$ as $x_{i}=\text{sign}(v_{i}^{\top}y)$. We use multiple hyperplanes and pick the best resulting solution.
@@ 4 Experiments
This section is organized as follows. First, we provide a comprehensive experimental evaluation of the OptGNN pipeline on several problems and datasets and compare it against solvers and neural baselines. We then describe a model ablation study, an out-of-distribution performance (OOD) study, and an experiment with our neural certification scheme.
4.1 Comparisons with classical algorithms
We report experimental measurements of the performance of the OptGNN approach on NP-Hard combinatorial optimization problems: Maximum Cut, Minimum Vertex Cover, and Maximum 3-SAT. We obtain results for several datasets and compare against greedy algorithms, local search, a state-of-the-art MIP solver (Gurobi), and various neural baselines. For Max-3-SAT and Min-Vertex-Cover we derive a quadratically penalized loss based on the SDP formulation of those problems. For details of the experimental setup see Appendix D.
@@ 4.1 Comparisons with classical algorithms
Walksat100 0.14±${}^{\pm\text{0.36}}$ (0.12) 0.36±${}^{\pm\text{0.52}}$ (0.25) 0.68±${}^{\pm\text{0.65}}$ (0.40)
Walksat1 0.94±${}^{\pm\text{0.92}}$ (0.01) 1.46±${}^{\pm\text{1.11}}$ (0.01) 1.97±${}^{\pm\text{1.28}}$ (0.01)
@@
Autograd 6.85±${}^{\pm\text{2.33}}$ (6.80) 7.52±${}^{\pm\text{2.38}}$ (6.75) 8.32±${}^{\pm\text{2.50}}$ (6.77)
Table 2: Average number of unsatisfied clauses for Max-3-SAT on random instances with $N=100$ variables and clause ratios $r=4.00,4.15,4.30$. Standard deviation is reported in superscript. In parentheses, we report the average time per instance on the test set in seconds. The number next to Walksat indicates the number of restarts.
Method RB200 RB500
CMP 1.031 ±${}^{\pm\text{0.006}}$ 1.015 ±${}^{\pm\text{0.004}}$
ErdősGNN 1.031 ±${}^{\pm\text{0.004}}$ 1.021 ±${}^{\pm\text{0.002}}$
Meta-EGN 1.028 ±${}^{\pm\text{0.005}}$ 1.016 ±${}^{\pm\text{0.002}}$
RUN-CSP 1.124 ±${}^{\pm\text{0.001}}$ 1.062 ±${}^{\pm\text{0.005}}$
OptGNN 1.012 ±${}^{\pm\text{0.006}}$ 1.013 ±${}^{\pm\text{0.003}}$
Greedy MVC 1.027 ±${}^{\pm\text{0.007}}$ 1.014 ±${}^{\pm\text{0.003}}$
Gurobi 10.0 (0.5s) 1.009 ±${}^{\pm\text{0.006}}$ 1.013 ±${}^{\pm\text{0.004}}$
Gurobi 10.0 (1s) 1.002 ±${}^{\pm\text{0.003}}$ 1.012 ±${}^{\pm\text{0.003}}$
Gurobi 10.0 (5s) 1.000 ±${}^{\pm\text{0.001}}$ 1.004 ±${}^{\pm\text{0.003}}$
Table 3: Average approximation ratio for vertex covers on forced RB instances, with standard deviations reported in superscript. Here, lower is better, and a ratio of 1.000 represents finding the minimum vertex cover.
Min-Vertex-Cover experiments. We evaluated OptGNN on forced RB instances, which are hard vertex cover instances from the RB random CSP model that contain hidden optimal solutions (Xu et al., 2007). We considered two forced RB distributions, RB200 and RB500, as specified in prior work (Wang & Li, 2023). The results are in Table 3, which also includes the performance of several neural and classical baselines as reported in (Wang & Li, 2023; Brusca et al., 2023). OptGNN consistently outperforms state-of-the-art unsupervised baselines on this task and is even able to match the performance of Gurobi with a 0.5s time limit.
Max-3-SAT experiment and ablation. The purpose of this experiment is twofold: to demonstrate the viability of OptGNN on the Max-3-SAT problem and to examine the role of overparameterization in OptGNN. We generate 3-SAT formulae on the fly with 100 variables and a random number of clauses in the $[400,430]$ interval, and train OptGNN for 100,000 iterations. We then test on instances with 100 variables and $\{400,415,430\}$ clauses. The results are in Table 2. We compare with WalkSAT, a classic local search algorithm for SAT, and ErdosGNN Karalias & Loukas (2020), a neural baseline trained in the same way. Furthermore, we compare with a simple baseline reported as “Autograd” that implements gradient descent directly on the Lagrangian. For details see D.1. OptGNN is able to outperform ErdősGNN consistently and improves significantly over Autograd, which supports the overparameterized message passing of OptGNN.
Max-Cut experiments. Table 6 presents a comparison between OptGNN, a greedy algorithm, and Gurobi for Max-Cut. OptGNN is competitive with Gurobi when run on a similar time budget and clearly outperforms greedy on all datasets. For results on more datasets see D.2. We generated Erdős-Renyi graphs $\mathcal{G}_{n,p}$ for $n\in[400,500]$ and $p=0.15$ on the fly and trained an OptGNN for 20k iterations. Table 4 shows test results on GSET benchmark instances (Benlic & Hao, 2013). Furthermore, we report the performance for several other baselines, including a greedy algorithm, the Goemans-Williamson SDP, and neural baselines. We can see that OptGNN consistently outperforms the SDP and the greedy algorithm, while also being competitive with the neural baselines. However, OptGNN is not able to outperform ANYCSP, which to the best of our knowledge achieves the current state-of-the-art results.
Method $N=800$ $N=1000$ $N=2000$ $N\geq 3000$
Greedy 411.44 359.11 737.00 774.25
SDP 245.44 $2$
RUNCSP 185.89 156.56 357.33 401.00
ECO-DQN 65.11 54.67 157.00 428.25
ECORD 8.67 8.78 39.22 187.75
ANYCSP 1.22 2.44 13.11 51.63
OptGNN 109.22 125.00 284.67 549.38
@@
Table 4: Comparison on GSET instances for Max-Cut against state of the art neural baselines. Numbers reported are the mean deviations from the best known Max-Cut values, reported as $f_{\text{best}}$ in table 2 of Benlic & Hao (2013). Lower is better.
Out of distribution generalization. We tested OptGNN’s ability to perform well on data distributions (for the same optimization problem) that it was not trained on. We test this using a collection of different datasets. The results can be seen in table 9 and subsection D.5. Overall, the results show that the model is capable of performing well even on datasets it has not been trained on.
Model ablation. We train modern GNN architectures from the literature with the same loss function and compare them against OptGNN. Please see D.3 for more details and results on multiple datasets for two different problems. Overall, OptGNN is the best performing model on both problems across all datasets.
@@
Figure 2: Experimental comparison of SDP versus OptGNN Dual Certificates on random graphs of 100 nodes for the Max-Cut problem. Our OptGNN certificates track closely with the SDP certificates while taking considerably less time to generate.
4.2 Experimental demonstration of neural certificates.
@@
In this section, we provide a simple experimental example of our neural certificate scheme on small synthetic instances. Deploying this scheme on Max-Cut on random graphs, we find this dual certificate to be remarkably tight. An example can be seen in figure 2. For $100$ node graphs with $1000$ edges our certificates deviate from the SDP certificate by about $20$ nodes but are dramatically faster to produce. The runtime is dominated by the feedforward of OptGNN which is $0.02$ seconds vs. the SDP solve time which is $0.5$ seconds on cvxpy. See C.2 for extensive discussion and additional results.

### v7

@@ 3.2 Message Passing for Max-CSPs
Let $\text{SDP}(\Lambda)$ be the optimal value of the SDP 1 on instance $\Lambda$. The approximation ratio for the Max-k-CSP problem achieved by the SDP 1 is $\min_{\Lambda\in\text{Max-k-CSP}}\frac{\text{OPT}(\Lambda)}{\text{SDP}(\Lambda)}$, where the minimization is taken over all instances $\Lambda$ with arity $k$. There is no polynomial time algorithm that can achieve a larger approximation ratio assuming the truth of the UGC raghavendra2008optimal. We construct our message passing algorithm as follows. First we introduce the definition of the vector form SDP and its associated quadratically penalized Lagrangian.
Definition (Quadratically Penalized Lagrangian).
Any standard form SDP $\Lambda$ can be expressed as the following vector form SDP for some matrix $V=[v_{1},v_{2},\dots,v_{N}]\in R^{N\times N}$.
$\min_{V\in\mathbb{R}^{N\times N}}\quad$ $\langle C,V^{T}V\rangle\quad\text{subject to:}\quad\langle A_{i},V^{T}V\rangle=b_{i}\quad\forall i\in[\mathcal{F}].$ (5)
For any SDP in vector form we define the $\rho$ quadratically penalized Lagrangian to be
$\mathcal{L}_{\rho}(V)\vcentcolon=\langle C,V^{T}V\rangle+\rho\sum_{i\in\mathcal{F}}\left(\langle A_{i},V^{T}V\rangle-b_{i}\right)^{2}.$ (6)
We show that gradient descent on this Lagrangian $\mathcal{L}_{\rho}(V)$ for the Max-CSP SDP 1 takes the form of a message-passing algorithm on the constraint graph that can provably converge to an approximate global optimum for the SDP (see algorithm 1). We see that gradient descent on $\mathcal{L}_{\rho}$ takes the form of a simultaneous message passing update on the constraint graph. See equation 60 and algorithm 3 for analytic form of the Max-CSP message passing update. See appendix A and B for analytic form of Min-Vertex-Cover and Max-3-SAT message passing updates. Our main theorem is then following.
@@
[Informal] Given a Max-k-CSP instance $\Lambda$ represented in space $\Phi=O(|\mathcal{P}|q^{k})$, there is a message passing Algorithm 3 on constraint graph $G_{\Lambda}$ with a per iteration update time of $O(\Phi)$ that computes in $O(\epsilon^{-4}\Phi^{4})$ iterations an $\epsilon$-approximate solution (solution satisfies constraints to error $\epsilon$ achieving objective value within $\epsilon$ of optimum) to SDP 1. For the formal theorem and proof see Theorem C.1.
3.3 OptGNN for Max-CSP
Next we define OptGNN for Max-CSP. See A and B for OptGNN for Vertex Cover and 3-SAT.
Definition (OptGNN for Max-CSP).
Let $\Lambda$ be a Max-CSP instance on a constraint graph $G_{\Lambda}$ with $N$ nodes. Let $U$ be an input matrix of dimension $r\times N$ for $N$ nodes with embedding dimension $r$. Let $\mathcal{L}_{\rho}$ be the penalized lagrangian loss defined as in equation 6 associated with the Max-CSP instance $\Lambda$. Let $M$ be the OptGNN weights which are a set of matrices $M\vcentcolon=\{M_{1},M_{2},...,M_{T}\}\in\mathbb{R}^{r\times 2r}$. Let $\text{LAYER}_{M_{i}}:\mathbb{R}^{r\times N}\rightarrow\mathbb{R}^{r\times N}$ be the function $\text{LAYER}_{M_{i}}(U)=M_{i}(\text{AGG}(U))$, where
$\text{AGG}:\mathbb{R}^{r\times N}\rightarrow\mathbb{R}^{2r\times N}$ is the aggregation function $\text{AGG}(U)\vcentcolon=[U,\nabla\mathcal{L}_{\rho}(U)]$. We define $\text{OptGNN}_{(M,\Lambda)}:\mathbb{R}^{r\times N}\rightarrow\mathbb{R}$ to be the function
$\text{OptGNN}_{(M,\Lambda)}(U)=\mathcal{L}_{\rho}\circ\text{LAYER}_{M_{T}}\circ\dots\circ\circ\text{LAYER}_{M_{1}}(U).$ (7)
The per iteration update time of OptGNN is $O(\Phi r^{\omega})$ where $\omega$ is the matrix multiplication exponent. We update the parameters of OptGNN by inputting the output of the final layer $\text{LAYER}_{M_{T}}$ into the Lagrangian $\mathcal{L}_{\rho}$ and backpropagate its derivatives. We emphasize the data is the instance $\Lambda$ and not the collection of vectors $U$ which can be chosen entirely at random. The form of the gradient $\nabla\mathcal{L}_{\rho}$ is a message passing algorithm over the nodes of the constraint graph $G_{\Lambda}$. Therefore, OptGNN is a message passing GNN over $G_{\Lambda}$ (see Definition Definition). This point is of practical importance as it is what informs out implementation of the OptGNN architecture. We then arrive at the following corollary.
Corollary 1 (Informal).
Given a Max-k-CSP instance $\Lambda$ represented in space $\Phi=O(|\mathcal{P}|q^{k})$, there is an OptGNN(M,Λ) with $T=O(\epsilon^{-4}\Phi^{4})$ layers, and embeddings of dimension $\Phi$ such that there is an instantiation of learnable parameters $M=\{M_{t}\}_{t\in[T]}$ that outputs a set of vectors $V$ satisfying the constraints of SDP 1 and approximating its optimum to error $\epsilon$. See formal statement 2
Moving on from our result on representing approximation algorithms, we ask whether OptGNN is learnable. That is to say, does OptGNN approximate the value of SDP 1 when given a polynomial amount of data? We provide a perturbation analysis to establish the polynomial sample complexity of PAC-learning OptGNN. The key idea is to bound the smoothness of the polynomial circuit AGG used in the OptGNN layer which is a cubic polynomial analogous to linear attention. We state the informal version below. For the formal version see Lemma E.5.
Lemma 3.1 (PAC learning).
Let $\mathcal{Q}$ be a dataset of Max-k-CSP instances over an alphabet of size $[q]$ with each instance represented in space $\Phi$. Here the dataset $\mathcal{D}\vcentcolon=\Lambda_{1},\Lambda_{2},...,\Lambda_{\Gamma}\sim\mathcal{D}$ is drawn i.i.d from a distribution over instances $\mathcal{D}$. Let $M$ be a set of parameters $M=\{M_{1},M_{2},...,M_{T}\}$ in a parameter space $\Theta$. Then for $T=O(\epsilon^{-4}\Phi^{4})$, for $\Gamma=O(\epsilon^{-4}\Phi^{6}\log^{4}(\delta^{-1}))$, let $\widehat{M}$ be the empirical loss minimizing weights for arbitrary choice of initial embeddings $U$ in a bounded norm ball. Then we have that OptGNN is $(\Gamma,\epsilon,\delta)$-PAC learnable. That is to say the empirical loss minimizer $\text{EMP}(\mathcal{Q})$ is within $\epsilon$ from the distributional loss with probability greater than $1-\delta$:
$\Pr\left[\left|\textup{\text{EMP}}(\mathcal{Q})-\mathbb{E}_{\Lambda\sim\mathcal{D}}[\text{OptGNN}_{(\widehat{M},\Lambda)}(U)]\right|\leq\epsilon\right]\geq 1-\delta.$
We note that this style of perturbation analysis is akin to the VC theory on neural networks adapted to our unsupervised setting. Although it’s insufficient to explain the empirical success of backprop, we believe our analysis sheds light on how architectures that capture gradient iterations can successfully generalize.
{wrapfigure}
@@ 3.3 OptGNN for Max-CSP
OptGNN in practice. Figure 1 summarizes the OptGNN pipeline for solving CO problems. OptGNN computes node embeddings $V\in\mathbb{R}^{N\times r}$ as per equation 7 which feeds into the loss $\mathcal{L}_{\rho}$. For pseudocode, please refer to the appendix (Max-Cut: algorithm 4 and general SDPs: algorithm 5). We use Adam (kingma2014adam) to update the parameter matrices $M$. Given a training distribution $\mathcal{D}$, the network is trained in a completely unsupervised fashion by minimizing $\mathbb{E}_{G\sim\mathcal{D}}[\mathcal{L}(\mathbf{v};G)]$ with a standard automatic differentiation package like PyTorch (paszke2019pytorch). A practical benefit of our approach is that users do not need to reimplement the network to handle each new problem. Users need only implement the appropriate loss, and our implementation uses automatic differentiation to compute the messages in the forward pass. At inference time, the output embeddings must be rounded to a discrete solution. To do this, we select a random hyperplane vector $y\in\mathbb{R}^{r}$, and for each node with embedding vector $v_{i}$, we calculate its discrete assignment $x_{i}\in\{-1,1\}$ as $x_{i}=\text{sign}(v_{i}^{\top}y)$. We use multiple hyperplanes and pick the best resulting solution.
@@
Obtaining neural certificates. We construct dual certificates of optimality from the output embeddings of OptGNN. The certificates provide a bound on the optimal solution of the SDP. The key idea is to estimate the dual variables of SDP 1. Since we use a quadratic penalty for constraints, the natural dual estimate is one step of the augmented method of Lagrange multipliers on the SDP solution which can be obtained straightforwardly from the primal problem variables $V$. We then analytically bound the error in satisfying the KKT conditions of SDP 1. See C.1 for derivations and extended discussion, and Figure 1 for an experimental demonstration.
Adaptation to Data vs. Theoretical Result. The theory result concerns worst case optimality where the embedding dimension is as large as $N$. In practice we deploy OptGNN with varying choices of the embedding dimension to obtain the best performance. For low dimensional embeddings, the corresponding low rank SDP objective is NP-hard to optimize. As such, training OptGNN is like training any other GNN for CO in that the weights adapt to the data. This is highlighted in our experiments where OptGNN generally outperforms SDP solvers.
@@ 4 Experiments
First, we provide a comprehensive experimental evaluation of the OptGNN pipeline on several problems and datasets and compare it against heuristics, solvers, and neural baselines. We then describe a model ablation study, an out-of-distribution performance (OOD) study, and an experiment with our neural certification scheme. We empirically test the performance of OptGNN on NP-Hard combinatorial optimization problems: Maximum Cut, Minimum Vertex Cover, and Maximum 3-SAT. We obtain results for several datasets and compare against greedy algorithms, local search, a state-of-the-art MIP solver (Gurobi), and various neural baselines. For Max-3-SAT and Min-Vertex-Cover we adopt the quadratically penalized Lagrangian loss of their respective SDP relaxations. For details of the setup see Appendix D.
@@
Table 2: Average number of unsatisfied clauses for Max-3-SAT on random instances with $N=100$ variables and clause ratios $r=4.00,4.15,4.30$. Standard deviation of the ratio over the test set is reported in superscript. In parentheses, we report the average time per instance on the test set in seconds.
@@
Walksat (100 restarts) 0.14±${}^{\pm\text{0.36}}$ (0.12) 0.36±${}^{\pm\text{0.52}}$ (0.25) 0.68±${}^{\pm\text{0.65}}$ (0.40)
Walksat (1 restart) 0.94±${}^{\pm\text{0.92}}$ (0.01) 1.46±${}^{\pm\text{1.11}}$ (0.01) 1.97±${}^{\pm\text{1.28}}$ (0.01)
Survey Propagation 3.32±${}^{\pm\text{0.81}}$ (0.001) 3.87±${}^{\pm\text{0.79}}$ (0.001) 3.94 ±${}^{\pm\text{0.93}}$ (0.001)
@@
Autograd SDP 6.85±${}^{\pm\text{2.33}}$ (6.80) 7.52±${}^{\pm\text{2.38}}$ (6.75) 8.32±${}^{\pm\text{2.50}}$ (6.77)
Low-Rank SDP ((wang2019low)) 12.38±${}^{\pm\text{1.06}}$(0.66) 13.32±${}^{\pm\text{1.09}}$ (0.67) 14.27±${}^{\pm\text{1.08}}$ (0.69)
@@
(a) Comparison on GSET Max-Cut instances against state-of-the-art neural baselines. Numbers reported are the mean (over the graphs in the test set) deviations from the best-known Max-Cut values, reported in benlic2013breakout.
(b) Average approximation ratio and standard deviation over the test set for vertex covers on forced RB instances. A ratio of 1.000 represents finding the minimum vertex cover.
Figure 1: Results for Max-Cut and Minimum Vertex Cover.
Min-Vertex-Cover experiments. We evaluated OptGNN on forced RB instances, which are hard vertex cover instances from the RB random CSP model that contain hidden optimal solutions (xu2007random). We use two distributions specified in prior work (wang2023unsupervised), RB200 and RB500. The results are in 1(b), which also includes the performance of several neural and classical baselines as reported in (wang2023unsupervised; brusca2023maximum). OptGNN consistently outperforms state-of-the-art unsupervised baselines on this task and is able to match the performance of Gurobi with a 0.5s time limit.
Max-3-SAT experiment and ablation. The purpose of this experiment is twofold: to demonstrate the viability of OptGNN on the Max-3-SAT problem and to examine the role of overparameterization in OptGNN. We generate 3-SAT formulae on the fly with 100 variables and a random number of clauses in the $[400,430]$ interval, and train OptGNN for 100,000 iterations. We then test on instances with 100 variables and $\{400,415,430\}$ clauses. The results are Table 2. We compare with WalkSAT, a classic local search algorithm for SAT, a low-rank SDP solver (wang2019satnet), Survey Propagation (braunstein2005survey), and ErdosGNN karalias2020erdos, a neural baseline trained in the same way. We also compare with a simple baseline reported as “Autograd” that directly employs gradient descent on the penalized Lagrangian using the autograd functionality of Pytorch. For details see D.2. OptGNN is able to outperform ErdősGNN consistently and improves significantly over Autograd, which supports the overparameterized message passing of OptGNN. OptGNN performs better than the low-rank SDP solver, though does not beat WalkSAT/Survey Prop. It is worth noting that the performance of OptGNN could likely be further improved without significant computational cost by applying a local search post-processing step to its solutions but we did not pursue this further in order to emphasize the simplicity of our approach.
Max-Cut experiments. Table 1 presents a comparison between OptGNN, a greedy algorithm, and Gurobi for Max-Cut. OptGNN clearly outperforms greedy on all datasets and is competitive with Gurobi when Gurobi is restricted to a similar runtime. For results on more datasets see subsection D.3. Following the experimental setup of ANYCSP tonshoff2022one, we also tested OptGNN on the GSET benchmark instances (benlic2013breakout). We trained an OptGNN for 20k iterations on generated Erdős-Renyi graphs $\mathcal{G}_{n,p}$ for $n\in[400,500]$ and $p=0.15$. 1(a) shows the results. We have included an additional low-rank SDP baseline to the results, while the rest of the baselines are reproduced as they appear in the original ANYCSP paper. These include state-of-the-art neural baselines, the Goemans-Williamson algorithm, and a greedy heuristic. We can see that OptGNN outperforms the SDP algorithms and the greedy algorithm, while also being competitive with the neural baselines. However, OptGNN does not manage to outperform ANYCSP, which to the best of our knowledge achieves the current state-of-the-art results for neural networks.
{wrapfigure}
@@
R0.5[Uncaptioned image] Experimental comparison of SDP versus OptGNN Dual Certificates on random graphs of 100 nodes for the Max-Cut problem. Our OptGNN certificates track closely with the SDP certificates while taking considerably less time to generate.
Out of distribution generalization. We test OptGNN’s ability to perform on data distributions (for the same optimization problem) that it’s not trained on. The results can be seen in table 7 and subsection D.6. The results show that the model is capable of performing well even on datasets it has not been trained on.
Model ablation. We train modern GNN architectures from the literature with the same loss function and compare them against OptGNN. Please see Appendix D.4 for more details and results on multiple datasets for two different problems. Overall, OptGNN is the best performing model on both problems across all datasets.
@@
Experimental demonstration of neural certificates. Next, we provide a simple experimental example of our neural certificate scheme on small synthetic instances. Deploying this scheme on Max-Cut on random graphs, we find this dual certificate to be remarkably tight. figure 1 shows an example. For $100$ node graphs with $1000$ edges our certificates deviate from the SDP certificate by about $20$ nodes but are dramatically faster to produce. The runtime is dominated by the feedforward of OptGNN which is $0.02$ seconds vs. the SDP solve time which is $0.5$ seconds on cvxpy. See C.1 for extensive discussion and additional results.